calculate how much energy is convered from chemical to electrical energy when 2C of charge flows from a 9V battery
1
Expert's answer
2020-05-18T10:07:34-0400

The energy convered from chemical to electrical energy


"E=I\cdot U\cdot t=q\cdot U=2\cdot9=18 J"